OpenStudy (anonymous):

I NEED HELP What is (-m)^-3 n if m=2 and n=-24 3 -3 4 -4

OpenStudy (anonymous):

(-m)^(-3)* n (-2)^(-3) *(-24) a^(-b)=1/a^(b) so (-2)^(-3)=1/(-2)^3 1/(-2)^3 * -24 -24/ -8= 3 So A(:

OpenStudy (shaik0124):

should i report u Ellolovee :) giving answer help not just giving answer

OpenStudy (anonymous):

Stfu , Um first off ! I showed them HOW to did it . I didnt just flat out give them the answer . so bye !

OpenStudy (shaik0124):

gl